# QWHA: Quantization-Aware Walsh-Hadamard Adaptation for Parameter-Efficient Fine-Tuning on Large Language Models

### v1

@@ D.1 Fine-tuning Hyperparameters
We follow the hyperparameter settings adapted from gao2024fourierft and deng2025cloq. Training is performed using the AdamW optimizer (loshchilov2017adamw). Table 8 reports the key settings, including minibatch size, weight decay, dropout ratio, learning rate scheduler, maximum sequence length, number of training epochs, warmup ratio, and the adapter scaling factor $\alpha$, which can be applied to adapters in Equation 2 through Equation 4. The scaling factors for low-rank and FT-based adapters are defined so that the $\ell_{2}$ norm of the gradients applied to their respective parameters is aligned. The learning rates for each combination of model, task, method, and bit-width are summarized in Table 9. We note that 128 sequences of length 2048, randomly sampled from the WikiText-2 (wiki2016merity) dataset, are used as a calibration set for quantization and adapter initialization, as these processes are robust to the choice of dataset (frantar2023gptq; zhang2024magrweightmagnitudereduction; deng2025cloq).
@@ D.2 Commonsense Question-Answering Benchmark
We present the detailed accuracy scores on the commonsense question answering (CSQA) benchmark in this section, while the averaged scores are reported in Section 4. The results for each of the seven individual tasks, are provided in Table 10 and Table D.2, corresponding to Table 4 and Table LABEL:table:eval_detail, respectively.

### v3

@@ D.1 Fine-tuning Hyperparameters
We follow the hyperparameter settings adapted from Gao et al. (2024b) and Deng et al. (2025). Training is performed using the AdamW optimizer (Loshchilov & Hutter, 2017). Table 8 reports the key settings, including minibatch size, weight decay, dropout ratio, learning rate scheduler, maximum sequence length, number of training epochs, warmup ratio, and the adapter scaling factor $\alpha$, which can be applied to adapters in Equation 2 through Equation 4. The scaling factors for low-rank and FT-based adapters are defined so that the $\ell_{2}$ norm of the gradients applied to their respective parameters is aligned. The learning rates for each combination of model, task, method, and bit-width are summarized in Table 9. We note that 128 sequences of length 2048, randomly sampled from the WikiText-2 (Merity et al., 2016) dataset, are used as a calibration set for quantization and adapter initialization, as these processes are robust to the choice of dataset (Frantar et al., 2023; Zhang et al., 2024; Deng et al., 2025).
@@ D.2 Commonsense Question-Answering Benchmark
We present the detailed accuracy scores on the commonsense question answering (CSQA) benchmark in this section, while the averaged scores are reported in Section 4. The results for each of the seven individual tasks, are provided in Table 10 and Table D.2, corresponding to Table 4 and Table 4.1, respectively.